# Progenitor diversity in the accreted stellar halos of Milky Way-like galaxies

### v1

@@ 3.2 Variation of diversity with galactocentric radius
Fig. 6 shows the main results of this paper: the separation of the total accreted stellar halo mass at a given radius into the fractional contributions from each progenitor. We call the diagrams in the top three rows of this figure ‘strata plots’ because of their passing resemblance to illustrations of geological strata. Contributions from a single progenitor over different radial bins (here ranging over 00 to $\sim 100\,\mathrm{kpc}$) are shaded with the same colour, and sorted vertically to place larger contributions relatively lower on the vertical scale of the diagram [^4] Unlike geological stratigraphy, the absolute vertical ‘height‘ of a layer on these diagrams has no physical significance. The different contributions to the accreted stellar halo are well-mixed in configuration space.. Naidu et al. (2020) used this diagram to illustrate their interpretation of observational constraints on the mix of stellar halo progenitors in the Milky Way; it has since been used in studies of the Auriga simulations by Fattahi et al. (2020) and Orkney et al. (2023). We will discuss the Milky Way data in the next section, after exploring similarities and differences between the strata plots from our three model sets.
@@
Figure 6: The strata plots and diversity profiles of each simulation over 1-100 kpc. The left panel is of BJ05 halos, the middle panel is of Aquarius halos and the right panel is of Auriga halos. For each simulation, the upper three panels are the strata plots; the lowest panel is the diversity profile of each halo, where the diversity is defined by Cooper et al. (2017). The strata plots show the fraction of accreted stellar halo mass contributed by individual progenitors at different galactic radii. Curves for different progenitors are stacked to form a strata look. The diversity profiles quantify the strata plots and show the number of significant progenitors on different radii (i.e. the more layers a strata plot has at the certain radius, the higher the $N_{sig}$ or the diversity is).
The strata plots provide further information on the origin of these trends. A smaller number of relatively massive progenitors dominate the Aquarius and Auriga strata, in contrast to most BJ05 halos. The lower diversity of Aquarius and Auriga overall is apparent, as is their trend of increasing diversity with radius. From the strata, we can further see that the rise in diversity with radius is usually driven by the relatively concentrated radial extent of debris from the dominant central progenitors, consistent with their early accretion and/or rapid in-spiral due to dynamical friction prior to disruption (e.g. Amorisco, 2017). The lower diversity in the outskirts of the Aquarius halos, which is not reflected in the majority of Auriga galaxies, is due to the contribution of a small number of comparably massive systems. This gives rise to apparent peaks in diversity at intermediate radii, where these progenitors do not dominate. This appears to be a systematic difference between the Aquarius and Auriga galaxies, the latter tending to have significantly more contribution from less massive progenitors in their outskirts (resembling the BJ05 galaxies, although with fewer such progenitors).
@@
The overall impression given by Fig. 5 is that range of diversity among the Aquarius halos is similar to that of the Auriga halos; however, the diversity profiles of individual halos in Fig. 6 show possible systematic differences between these two sets of simulations.
Since Aquarius and Auriga are both cosmological $N$-body simulations of similar Milky Way-like systems, it seems unlikely that these differences arise only from halo-to-halo variations in progenitor mass or accretion time. Nevertheless, the star formation histories, and consequently the stellar masses, of the progenitors may be quite different in the two sets of models. For example, if a given model forms systematically more stars in progenitors that are accreted at early times, that may be reflected in the diversity profiles. Differences between particle tagging (Aquarius) and more self-consistent hydrodynamics (Auriga) may also be relevant. These factors were discussed by Cooper et al. (2017), who showed (for a single example, Aq-C) that differences in star formation histories predicted by semi-analytic and hydrodynamical models applied to the same initial conditions were likely to dominate differences in the diversity profile, while the limitations of particle tagging contributed in more subtle ways. The absence of a massive central disk potential in Aquarius may be particularly important, because such a potential could strongly affect the orbits of the most massive progenitors (for example, progenitors may be disrupted earlier, and their orbits may be circularized in the plane of the disk). In that regard, we note that the BJ05 simulations, which include a (rigid) disk potential, do not recover behavior resembling the Auriga galaxies in the central regions of the halo, although this may be due to other differences between BJ05 and Auriga.
Also notable in the Auriga strata plots are the ‘spikes’ that appear in the outer regions of some halos. These arise from unbound material in the vicinity of surviving satellites that are undergoing tidal disruption. Since these stars are unbound from their original satellites, our definition counts them as part of the accreted halo. They appear to have a disproportionate significance in these plots at large radii, because the total mass of the ‘smooth’ accreted halo at those radii is so low (the strata plots show only relative contributions). At face value, similar features should be expected in Aquarius, but we do not see them. There are several possible explanations for why they are more prominent in Auriga: Auriga satellites may survive for longer in the final phases of tidal stripping due to additional binding by their stellar component; the satellites themselves may have more extensive stellar halos in Auriga, leading to earlier low-level stellar mass loss; and strong interactions with the disk in Auriga may produce streams with more unbound stars piling up near to their former parent satellite at apocenter.
@@ 3.3 Comparison to the Milky Way
In this section, we compare the simulations above to current constraints on the composition of the Milky Way’s accreted stellar halo. Naidu et al. (2020) present a census of likely contributions to the MW stellar halo in the range $8-45$ kpc, based on the literature and their own measurements from the H3 spectroscopic survey. This radial range corresponds to Figure 19 of Naidu et al. (2020), which directly inspired our ‘strata plot’ presentation. Naidu et al. (2020) attempt to correct their spectroscopic sample for various sources of incompleteness in their selection function, but not for possible bias due to limited sky coverage; their data are restricted to a high latitude footprint and, within that footprint, to heights above the disk plane $|Z|>2\,\mathrm{kpc}$. It is possible that some progenitors may make significant contributions to the volume that Naidu et al. do not probe. For example, Gómez et al. (2017) discuss examples from simulations in which much of the debris from massive progenitors is deposited in the plane of the disk.
To simplify and generalize the presentation here, particularly with regard to comparisons between the different models, we do not impose these geometric limits on the simulations. This simplification should be kept in mind when assessing the differences between the simulations and the results of Naidu et al. [^5] Kizhuprakkat et al. (2024) constructed mock catalogs for the DESI Milky Way Survey, which has a qualitatively similar footprint to H3, based on the L3 Auriga simulations. They find that the 10 most massive accreted progenitors are visible in such a footprint, regardless of the azimuthal position of the Sun in the simulated galaxy. Moreover, their relative contributions in the DESI footprint closely match their relative contributions to the halo overall..
Naidu et al. (2020) present their interpretation in terms of mass fractions rather than absolute masses. Our definition of $N_{\mathrm{sig}}$ in equation 1 can be re-written in terms of mass fractions:
@@
Figure 7: The diversity profile of accreted stellar halos. From top to bottom, panels show the BJ05 halos; the Aquarius halos; and the Auriga halos, split into three groups according to the virial mass of each galaxy. Thicker and darker grey lines indicate relatively higher accreted stellar halo mass in each group. The shorter black line, repeated in every panel, shows the Milky Way diversity profile we have computed based on the data in Naidu et al. (2020).
@@
In Fig. 7, we compare the diversity profiles of selected examples from the models to that of the Milky Way, which we computed based on the information given in Figure 19 of Naidu et al. (2020). We carry out this comparison at face value, without examining the assumptions underlying the interpretation of Naidu et al. (2020). We have excluded from our comparison the fraction of mass they associate with in-situ stars. Surprisingly, we find the implied diversity of the Milky Way is low compared to almost all of the models. In contrast to most simulated halos, 75% of the Milky Way halo appears to be contributed by only two massive progenitors, GS/E and Sagittarius, across the entire range we study.  [^6] We note that some proposed progenitors are not currently identified in the H3 sample (e.g. Kruijssen et al., 2018; Horta et al., 2020; Kruijssen et al., 2020).
@@
In Fig. 8, we show strata plots in the range $8-45$ kpc for a subset of simulated halos from the three models that we judge (by eye) to have diversity profiles most similar to that of the Milky Way (in Fig. 7). We again compare to (a slightly modified version of) Figure 19 from Naidu et al. (2020), shown in the central panel [^7] Note that the radii in this figure are galactocentric, not heliocentric. In this reconstruction, the contribution of the Saggitarius stream at radii $<10$ kpc is on the opposite side of the Galactic centre from the Sun..
@@
Figure 8: A subset of strata plots for halos that have low diversity and ‘flat’ diversity profiles to $\sim 40$ kpc, resembling current inferences for the Milky Way (central panel, see text). The majority of halos selected in this way experience substantial (low mass ratio) mergers close to the present day (see Fig. 9).
@@
These ‘best-matching’ halos are among the least typical in each model set. As shown in Fig. 5, the most massive progenitor of Aq-F contributes around 98% of the total accreted stellar halo mass. This galaxy has undergone a recent major merger and hence is unlikely to have an undisturbed disk resembling the Milky Way. In projection its halo is dominated by a prominent system of ‘shells’ (see Cooper et al., 2011). Similarly, Fattahi et al. (2019) exclude Au11 and Au20 from their study on the basis that they are still undergoing merger events at $z=0$. Monachesi et al. (2019) exclude Au11 and Au01, which they find to have a nearly equal mass companion nearby. These four halos could therefore be said to be among the least ‘Milky Way-like’ of the Auriga set, due to their recent or ongoing merging. In contrast to these examples, the Milky Way seems to have atypically low diversity, despite (apparently) not having undergone a recent major merger.
@@
Figure 9: As Figure 8, but here with strata colored according to the infall time of each progenitor. The estimated infall times of GS/E and Sgr are taken from Xiang & Rix (2022) and Hughes et al. (2019) respectively. We colour the other Milky Way progenitors blue to indicate that their infall times are not yet well constrained.
@@
The discrepancy between the Milky Way and the simulations is even clearer in Fig. 9, where we colour-code the strata according to the time of infall of the progenitor (an upper limit to the lookback time at which the stars were stripped into the halo). The current consensus is that GS/E originated in an ‘early’ merger ($\gtrsim 11\,\mathrm{Gyr}$ ago, e.g. Xiang & Rix, 2022). Although such early mergers do occur in the models (e.g. in Au23), they are rarely dominant contributors to the present-day stellar halo. When a single progenitor dominates the present-day halo, that progenitor typically arrives relatively late ($\lesssim 9$ Gyr ago), as in Au20, Au11, Au01, Aq-F and BJ08.
@@
Fattahi et al. (2019) searched the Auriga halos for features that resemble the observed spatial, chemical and kinematic properties of GS/E (see also Orkney et al., 2023). None of the Auriga halos in Fig. 8 are GS/E analogues according to their criteria. Conversely, none of the GS/E-analogue Auriga halos identified by Fattahi et al. (2019) have diversity profiles resembling that of the Milky Way. Superficially ‘GS/E-like’ (early-accreted, centrally concentrated, high mass fraction) features in the strata plots of halos with somewhat Milky Way-like diversity profiles (for example, Au08) are not GS/E-like according to the criteria of Fattahi et al. (2019). Furthermore, we do not see any notable similarity among the strata plots or diversity profiles of the Fattahi et al. (2019) GS/E analogs.
@@
Essentially the same considerations apply in the case of Sgr; the only difference is a somewhat later infall time and a more circular orbit, leading to more deposition in the outer halo. The infall time of Sgr is thought to be $\sim 9\,\mathrm{Gyr}$ ago (e.g. Hughes et al., 2019), and the progenitor has not yet been completely disrupted. In a halo dominated by a single massive GS/E-like progenitor, the late arrival of a comparably massive Sgr-like progenitor could produce a diversity distribution that resembles the picture sketched for the Milky Way. Might the Milky Way have been more consistent with the models before Sagittarius arrived? At face value, the answer is no – the models we have examined suggest that a late-type galaxy with an accreted stellar halo dominated by a single GS/E analog would be even more unusual. The state after the accretion of a late-arriving Sgr analogue is arguably closer to typical.
@@
Therefore, taking the models at face value, the Milky Way appears unusual. We do not find any halos out of the 47 Milky Way analogues from these three different models that resemble the combination of apparently low diversity and early assembly inferred for the Milky Way. Among those simulated galaxies that do not experience a major merger, few (essentially none in our sample) have two such dominant progenitors as GS/E and Sgr arriving so early. Moreover, the latest-arriving progenitor is usually the most massive. The only clear exception is Aq-A, an atypically early-forming dark matter halo, in which a relatively early merger (the most massive) dominates the accreted stellar halo alongside two even earlier mergers of comparable mass. Removing one of the two more ancient Aq-A progenitors (or interpreting observations of their debris as a single event) would produce a somewhat Milky Way-like result [^8] These mergers in Aq-A also involve a substantial fraction of stars on a wide, somewhat Sgr-like orbit; see Cooper et al. (2010)..
@@
The statistics of this comparison are quite poor; the Milky Way is known to be an atypical galaxy for its mass in several respects (for example, the presence of the LMC (e.g. Boylan-Kolchin et al., 2011), and its apparently relatively efficient in situ star formation (Licquia & Newman, 2015; Licquia et al., 2016; Boardman et al., 2020). If being an outlier in these respects correlates with accreted stellar halo diversity [^9] A reasonable speculation, because it almost certainly reflects aspects of the halo growth history., it is not surprising that we do not find a good match in a small sample of models based on this single point of comparison. This is compounded by the fact that the models span a wide range of virial mass, and by uncertainty on the true virial mass of the Milky Way. Our sample may include only a handful of galaxies (or none) that are appropriate mass-analogues of the Milky Way.
@@
Alternatively, the low diversity inferred for the Milky Way may be an underestimate. Other debris systems that have not been accounted for may contribute at a significant level. Perhaps the most likely reason for this underestimate would be that multiple systems have been classified with a single hypothetical progenitor in the current interpretation. This seems more plausible for GS/E, the full extent and classification of which remains poorly understood (see Donlon et al., 2022, 2023; Sharpe et al., 2024; Carrillo et al., 2024). It seems less plausible for Sagittarius.
@@
To illustrate this point, the upper two panels of Fig. 10 we show strata plots for Au03 and Au09, two halos with higher diversity than the Milky Way, but a similar accreted stellar halo assembly history – that is, one or more progenitors of dominant, GS/E-like centrally concentrated debris with infall time 10-11 Gyr ago, and one or more progenitors of less concentrated Sgr-like debris with infall time 9-10 Gyr ago. Au09 also has an accreted stellar halo mass and metallicity gradient similar to the Milky Way (Monachesi et al., 2019). In a comparable range of radius, we find these halos have strata much more closely resembling those of Naidu et al. (2020) in Fig. 9. An interpretation of the Milky Way data in which ”GS/E” or ”Sgr” debris was attributed to more than one progenitor with similar masses and infall times would be comparable to these models.
@@
As above, we can consider the effect of stochastic accretion on this comparison; the lowest panel of Fig. 10 shows the Au08 strata after removing the contribution of the dominant recent accretion event ($\lesssim 9$ Gyr ago) removed. The remaining strata are the closest match to the Milky Way interpretation of Naidu et al. (2020) that we can identify in our models, the only difference being that both the ‘GS/E’ analogue and the ‘Sgr’ analogue each comprise two systems with similar radial distribution.
@@
Fig. 10 is plotted over a larger range of radius; although Naidu et al. (2020) probe distances up to $\sim 40$ kpc with sufficient density to attempt estimates of the mass fractions of different halo structures, this is still a relatively small scale in terms of the accreted stellar halo overall. Although the strata of the three examples are similar in the region probed by H3, Au09 has many more strata (higher diversity) than Au03 and Au08 at radii $<8$ kpc and $>40$ kpc. The outer region of Au08 is dominated by debris associated with a very recently accreted satellite (the black “hills” around $\sim 100$ kpc). Most simulated diversity profiles rise at larger radii in Auriga (although not in Aquarius). Evidence for or against the increasing diversity of the Milky Way halo at $60$-$100$ kpc would therefore provide annother useful point of comparison with models.

### v2

@@ 3.2 Variation of diversity with galactocentric radius
Figure 6 shows the main results of this paper: the separation of the total accreted stellar halo mass at a given radius into the fractional contributions from each progenitor. We call the diagrams in the top three rows of this figure ‘strata plots’ because of their passing resemblance to illustrations of geological strata. Contributions from a single progenitor over different radial bins (here ranging over 00 to $\sim 100\,\mathrm{kpc}$) are shaded with the same color, and they have been sorted vertically to place larger contributions relatively lower on the vertical scale of the diagram [^4] Unlike geological stratigraphy, the absolute vertical ‘height’ of a layer on these diagrams has no physical significance. The different contributions to the accreted stellar halo are well-mixed in configuration space.. Naidu et al. (2020) used this diagram to illustrate their interpretation of observational constraints on the mix of stellar halo progenitors in the Milky Way; it has since been used in studies of the Auriga simulations by Fattahi et al. (2020) and Orkney et al. (2023). We will discuss the Milky Way data in the next section, after exploring similarities and differences between the strata plots from our three model sets.
@@
Figure 6: The strata plots and diversity profiles of each simulation over 1-100 kpc. The left panel consists of BJ05 halos, the middle panel shows Aquarius halos, and the right panel displays Auriga halos. For each simulation, the upper three panels are the strata plots; the lowest panel is the diversity profile of each halo, where the diversity is defined by Cooper et al. (2017). The strata plots show the fraction of accreted stellar halo mass contributed by individual progenitors at different galactic radii. Curves for different progenitors are stacked to form a strata look. The diversity profiles quantify the strata plots and show the number of significant progenitors on different radii (i.e. the more layers a strata plot has at the certain radius, the higher the $N_{sig}$ or the diversity is).
The strata plots provide further information on the origin of these trends. A smaller number of relatively massive progenitors dominate the Aquarius and Auriga strata, in contrast to most BJ05 halos. The lower diversity of Aquarius and Auriga overall is apparent, as is their trend of increasing diversity with radius. From the strata, we can further see that the rise in diversity with radius is usually driven by the relatively concentrated radial extent of debris from the dominant central progenitors, consistent with their early accretion and/or rapid inspiral due to dynamical friction prior to disruption (e.g. Amorisco, 2017a). The lower diversity in the outskirts of the Aquarius halos, which is not reflected in the majority of Auriga galaxies, is due to the contribution of a small number of comparably massive systems. This gives rise to apparent peaks in diversity at intermediate radii, where these progenitors do not dominate. This appears to be a systematic difference between the Aquarius and Auriga galaxies, with the latter tending to have significantly more contribution from less massive progenitors in their outskirts (resembling the BJ05 galaxies, although with fewer such progenitors).
@@
The overall impression given by Figure 5 is that range of diversity among the Aquarius halos is similar to that of the Auriga halos; however, the diversity profiles of individual halos in Figure 6 show possible systematic differences between these two sets of simulations.
Since Aquarius and Auriga are both cosmological $N$-body simulations of similar Milky Way–like systems, it seems unlikely that these differences arise only from halo-to-halo variations in progenitor mass or accretion time. Nevertheless, the star formation histories, and consequently the stellar masses, of the progenitors may be quite different in the two sets of models. For example, if a given model forms systematically more stars in progenitors that are accreted at early times, that may be reflected in the diversity profiles. Differences between particle tagging (Aquarius) and more self-consistent hydrodynamics (Auriga) may also be relevant. These factors were discussed by Cooper et al. (2017), who showed (for a single example, Aq-C) that differences in star formation histories predicted by semi-analytic and hydrodynamical models applied to the same initial conditions were likely to dominate differences in the diversity profile, while the limitations of particle tagging contributed in more subtle ways. The absence of a massive central disk potential in Aquarius may be particularly important, because such a potential could strongly affect the orbits of the most massive progenitors (for example, progenitors may be disrupted earlier, and their orbits may be circularized in the plane of the disk). In that regard, we note that the BJ05 simulations, which include a (rigid) disk potential, do not recover behavior resembling that of the Auriga galaxies in the central regions of the halo, although this may be due to other differences between BJ05 and Auriga.
@@ 3.3 Comparison to the Milky Way
In this section, we compare the simulations above to current constraints on the composition of the Milky Way’s accreted stellar halo. Naidu et al. (2020) present a census of likely contributions to the MW stellar halo in the range 8–45 kpc, based on the literature and their own measurements from the H3 spectroscopic survey. This radial range corresponds to Figure 19 of Naidu et al. (2020), which directly inspired our ‘strata plot’ presentation. Naidu et al. (2020) attempt to correct their spectroscopic sample for various sources of incompleteness in their selection function, but not for possible bias due to limited sky coverage; their data are restricted to a high-latitude footprint—and, within that footprint, to heights above the disk plane $|Z|>2\,\mathrm{kpc}$. It is possible that some progenitors may make significant contributions to the volume that Naidu et al. do not probe. For example, Gómez et al. (2017) discuss examples from simulations in which much of the debris from massive progenitors is deposited in the plane of the disk.
To simplify and generalize the presentation here, in particular with regard to comparisons between the different models, we do not impose these geometric limits on the simulations. This simplification should be kept in mind when assessing the differences between the simulations and the results of Naidu et al. [^5] Kizhuprakkat et al. (2024) constructed mock catalogs for the DESI Milky Way Survey, which has a qualitatively similar footprint to H3, based on the L3 Auriga simulations. They find that the 10 most massive accreted progenitors are visible in such a footprint, regardless of the azimuthal position of the Sun in the simulated galaxy. Moreover, their relative contributions in the DESI footprint closely match their relative contributions to the halo overall.
Naidu et al. (2020) present their interpretation in terms of mass fractions rather than absolute masses. Our definition of $N_{\mathrm{sig}}$ in equation 1 can be rewritten in terms of mass fractions:
@@
Figure 7: The diversity profile of accreted stellar halos. From top to bottom, panels show the BJ05 halos, the Aquarius halos, and the Auriga halos, split into three groups according to the virial mass of each galaxy. Thicker and darker gray lines indicate relatively higher accreted stellar halo mass in each group. The shorter black line, repeated in every panel, shows the Milky Way diversity profile we have computed based on the data in Naidu et al. (2020).
@@
In Figure 7, we compare the diversity profiles of selected examples from the models to that of the Milky Way, which we computed based on the information given in Figure 19 of Naidu et al. (2020). We carry out this comparison at face value, without examining the assumptions underlying the interpretation of Naidu et al. (2020). We have excluded from our comparison the fraction of mass they associate with in situ stars. Surprisingly, we find the implied diversity of the Milky Way is low compared to almost all of the models. In contrast to most simulated halos, 75% of the Milky Way halo appears to be contributed by only two massive progenitors, GS/E and Sagittarius, across the radial range we study [^6] Some potentially significant debris features that have been proposed are not currently identified in the H3 sample, notably those in the inner region of the galaxy (e.g. Kruijssen et al., 2018; Horta et al., 2020; Kruijssen et al., 2020). It is not yet clear if these are separate features and distinct from the in situ halo in the inner galaxy (see, e.g., section 4.3.2 of Deason & Belokurov, 2024). Horta et al. (2024) find analogous features in simulations are associated with progenitors that deposit the bulk of their mass within $<10$ kpc. We do not expect the absence of these features in H3 to have a strong effect on our comparison with Naidu et al. (2020). If (for example) some fraction of the in situ Naidu et al. (2020) is really accreted debris associated with an inner-Galaxy structure, this would give only a small increase in diversity up to $\sim 15$ kpc..
@@
In Figure 8, we show strata plots in the range $8-45$ kpc for a subset of simulated halos from the three models that we judge (by eye) to have diversity profiles most similar to that of the Milky Way (in Figure 7). We again compare to (a slightly modified version of) Figure 19 from Naidu et al. (2020), shown in the central panel [^7] Note that the radii in this figure are galactocentric, not heliocentric. In this reconstruction, the contribution of the Sagittarius stream at radii $<10$ kpc is on the opposite side of the Galactic center from the Sun..
@@
Figure 8: A subset of strata plots for halos that have low diversity and ‘flat’ diversity profiles to $\sim 40$ kpc, resembling current inferences for the Milky Way (central panel, see text). The majority of halos selected in this way experience substantial (low mass ratio) mergers close to the present day (see Figure 9).
@@
These ‘best-matching’ halos are among the least typical in each model set. As shown in Figure 5, the most massive progenitor of Aq-F contributes around 98% of the total accreted stellar halo mass. This galaxy has undergone a recent major merger and hence is unlikely to have an undisturbed disk resembling the Milky Way. In projection, its halo is dominated by a prominent system of ‘shells’ (see Cooper et al., 2011). Similarly, Fattahi et al. (2019) exclude Au11 and Au20 from their study on the basis that they are still undergoing merger events at $z=0$. Monachesi et al. (2019) exclude Au11 and Au01, which they find to have a nearly equal-mass companion nearby. These four halos could therefore be said to be among the least ‘Milky Way—like’ of the Auriga set, due to their recent or ongoing merging. In contrast to these examples, the Milky Way seems to have atypically low diversity, despite (apparently) not having undergone a recent major merger.
@@
Figure 9: As Figure 8, but here with strata colored according to the infall (lookback) time of each progenitor. The estimated infall times of GS/E and Sgr are taken from Xiang & Rix (2022) and Hughes et al. (2019), respectively. We color the other Milky Way progenitors blue to indicate that their infall times are not yet well constrained.
@@
The discrepancy between the Milky Way and the simulations is even clearer in Figure 9, where we color code the strata according to the time of infall of the progenitor (an upper limit to the lookback time at which the stars were stripped into the halo). The current consensus is that GS/E originated in an ‘early’ merger ($\gtrsim 11\,\mathrm{Gyr}$ ago, e.g., Xiang & Rix, 2022). Although such early mergers do occur in the models (e.g., in Au23), they are rarely dominant contributors to the present-day stellar halo. When a single progenitor dominates the present-day halo, that progenitor typically arrives relatively late ($\lesssim 9$ Gyr ago), as in Au20, Au11, Au01, Aq-F, and BJ08.
@@
Fattahi et al. (2019) searched the Auriga halos for features that resemble the observed spatial, chemical, and kinematic properties of GS/E (see also Orkney et al., 2023). None of the Auriga halos in Figure 8 are GS/E analogs according to their criteria. Conversely, none of the GS/E-analog Auriga halos identified by Fattahi et al. (2019) have diversity profiles resembling that of the Milky Way. Superficially ‘GS/E-like’ (early-accreted, centrally concentrated, high mass fraction) features in the strata plots of halos with somewhat Milky Way–like diversity profiles (for example, Au08) are not GS/E-like according to the criteria of Fattahi et al. (2019). Furthermore, we do not see any notable similarity among the strata plots or diversity profiles of the Fattahi et al. (2019) GS/E analogs.
@@
Essentially the same considerations apply in the case of Sgr; the only difference is a somewhat later infall time and a more circular orbit, leading to more deposition in the outer halo. The infall time of Sgr is thought to be $\sim 9\,\mathrm{Gyr}$ ago (e.g. Hughes et al., 2019), and the progenitor has not yet been completely disrupted. In a halo dominated by a single massive GS/E-like progenitor, the late arrival of a comparably massive Sgr-like progenitor could produce a diversity distribution that resembles the picture sketched for the Milky Way. Might the Milky Way have been more consistent with the models before Sagittarius arrived? At face value, the answer is no – the models we have examined suggest that a late-type galaxy with an accreted stellar halo dominated by a single GS/E analog would be even more unusual. The state after the accretion of a late-arriving Sgr analog is arguably closer to typical.
@@
Therefore, taking the models at face value, the Milky Way appears unusual. We do not find any halos out of the 47 Milky Way analogs from these three different models that resemble the combination of apparently low diversity and early assembly inferred for the Milky Way. Among those simulated galaxies that do not experience a major merger, few (essentially none in our sample) have two such dominant progenitors as GS/E and Sgr arriving so early. Moreover, the latest-arriving progenitor is usually the most massive. The only clear exception is Aq-A, an atypically early-forming dark matter halo, in which a relatively early merger (the most massive) dominates the accreted stellar halo alongside two even earlier mergers of comparable mass. Removing one of the two more ancient Aq-A progenitors (or interpreting observations of their debris as a single event) would produce a somewhat Milky Way–like result [^8] These mergers in Aq-A also involve a substantial fraction of stars on a wide, somewhat Sgr-like orbit; see Cooper et al. (2010).. This apparent inconsistency with simulations also motivates consideration of more recent accretion times for the dominant Milky Way halo structures. For example, Donlon et al. (2019, 2020, 2022), Donlon & Newberg (2023) and Donlon et al. (2024) examine an alternative scenario in which the GS/E progenitor (or some fraction of the debris associated with GS/E) merged with the Milky Way in the last 2-3 Gyr. Although this is even more recent than most of the dominant progenitors in the models we consider, a GS/E-like accretion in the last $\sim 5$ Gyr would be consistent with the overall trend in the simulations (see also Folsom et al., 2024).
@@
The statistics of this comparison are quite poor; the Milky Way is known to be an atypical galaxy for its mass in several respects (for example, the presence of the LMC (e.g. Boylan-Kolchin et al., 2010; Liu et al., 2011; Shao et al., 2018; Engler et al., 2021), and its apparently relatively efficient in situ star formation (Licquia & Newman, 2015; Licquia et al., 2016; Boardman et al., 2020)). If being an outlier in these respects correlates with accreted stellar halo diversity, [^9] A reasonable speculation, because it almost certainly reflects aspects of the halo growth history. it is not surprising that we do not find a good match in a small sample of models based on this single point of comparison. This is compounded by the fact that the models span a wide range of virial mass, and by uncertainty on the true virial mass of the Milky Way. Our sample may include only a handful of galaxies (or none) that are appropriate mass analogs of the Milky Way.
@@
Alternatively, the low diversity inferred for the Milky Way may be an underestimate. Other debris systems that have not been accounted for may contribute at a significant level. Perhaps the most likely reason for this underestimate would be that multiple systems have been identified with a single hypothetical progenitor in the current interpretation of the Milky Way data. This has been proposed for both GS/E (e.g. Donlon et al., 2022, 2024; Sharpe et al., 2024; Carrillo et al., 2024) and Sgr (e.g. Newby et al., 2015; Davies et al., 2024). Folsom et al. (2024) posit that $\sim 1/3$ of GS/E-like debris systems in the TNG50 simulations had multiple progenitors.
@@
To illustrate this point, the upper two panels of Figure 10 show strata plots for Au03 and Au09, two halos with higher diversity than the Milky Way but a similar accreted stellar halo assembly history – that is, one or more progenitors of dominant, GS/E-like centrally concentrated debris with infall time 10-11 Gyr ago and one or more progenitors of less concentrated Sgr-like debris with infall time 9-10 Gyr ago. Au09 also has an accreted stellar halo mass and a metallicity gradient similar to those of the Milky Way (Monachesi et al., 2019). In a comparable range of radius, we find these halos have strata much more closely resembling those of Naidu et al. (2020) in Figure 9. An interpretation of the Milky Way data in which ‘GS/E’ or ‘Sgr’ debris was attributed to more than one progenitor with similar masses and infall times would be comparable to these models.
@@
As above, we can consider the effect of stochastic accretion on this comparison; the lowest panel of Figure 10 shows the Au08 strata after removing the contribution of the dominant recent accretion event ($\lesssim 9$ Gyr ago) removed. The remaining strata are the closest match to the Milky Way interpretation of Naidu et al. (2020) that we can identify in our models—the only difference being that both the ‘GS/E’ analog and the ‘Sgr’ analog each comprise two systems with similar radial distribution.
@@
Figure 10 is plotted over a larger range of radius; although Naidu et al. (2020) probe distances up to $\sim 40$ kpc with sufficient density to attempt estimates of the mass fractions of different halo structures, this is still a relatively small scale in terms of the accreted stellar halo overall. Although the strata of the three examples are similar in the region probed by H3, Au09 has many more strata (higher diversity) than Au03 and Au08 at radii $<8$ kpc and $>40$ kpc. The outer region of Au08 is dominated by debris associated with a very recently accreted satellite (the black ‘hills’ around $\sim 100$ kpc). Most simulated diversity profiles rise at larger radii in Auriga (although not in Aquarius). Evidence for or against the increasing diversity of the Milky Way halo at $60$-$100$ kpc would therefore provide another useful point of comparison with models.